# Surprisingly fragile: Assessing and addressing prompt instability in multimodal foundation models

### v1

@@ Abstract
Introduction
While text-only large language models (LLMs) are powerful in their own right, multimodal foundation models offer even more transformative capabilities such as image captioning and audio signal processing (li2024). Generalist models such as Unified-IO are trained to process a range of modalities including image, video, audio, and depth maps, often using an LLM as a combined encoder for all input types  (lu2022). MFMs have the potential for major impact in a variety of downstream applications such as autonomous navigation (cui2024) and medical image analysis (moor2023), and many public-facing LLMs such as ChatGPT are increasingly incorporating multimodal modeling into the user experience to maximize their impact.
Despite their power, these models often show surprising fragility in the face of data that deviates slightly from the expected format: a model may respond very differently to a prompt that has a slight difference in phrasing (newman2021). We show an example from a popular video QA dataset (iVQA) in Table 1, in which a person fixes a car engine and the model is asked to respond to the prompt “What mechanism is the person working on?”. In response to prompt perturbations, the Unified-IO model with no additional training responds with a variety of incorrect responses, while the model that has been trained on prompt perturbations (“Joint similarity”; see Prompt augmentation) responds correctly regardless of phrasing.
@@ Introduction
We investigate the relative stability of current state-of-the-art MFMs on several standard multimodal benchmark datasets and find the following:
@@
  • Modern MFMs show a severe drop in performance in response to perturbed prompts on standard QA datasets (Testing on perturbed data).
@@
  • Re-training generalist multimodal models on a wider variety of perturbed prompts, regardless of how the prompts are selected (Prompt augmentation), improves performance and reduces instability on other prompt perturbations, more than simply re-training on the original prompts from the task data (Training on perturbed data). The perturbation-trained models show consistent improvement when the prompts target modality-specific information, e.g. prompts in the image modality that reference material properties (Error analysis).
@@
Our work provides an automated method to help align MFMs to human expectations, such as consistency in response to input variation (zhao2024).
Related work
Multimodal foundation models
@@
Recent work has unified text-based foundation models with multimodal processing to yield multimodal foundation models that can respond to the combination of text and multimodal input (li2024). The majority of such models, including Unified-IO and OFASys, first encode the text and multimodal data to a shared representation space, then decode the combined representation into the desired modality, often text (e.g. image QA requires converting image + text to text). Some systems such as Next-GPT learn to perform any-to-any modality generation by training modality-specific decoders (wu2023), which works reasonably well for more common modalities such as images. For breadth, systems including ImageBind (girdhar2023) and OneLLM (han2023) use a universal encoder and learn from a wide variety of data including depth maps and point clouds, which may help the model’s internal LLM achieve better cross-modality understanding. To further improve performance, parameter-efficient methods such as prompt learning have helped to solve problems with missing data (lee2023) and cross-modality alignment (khattak2023). We identify a possible gap in the research space, i.e. model-agnostic methods to improve robustness, and test a variety of prompt perturbation methods to determine impact on trained model performance.
Model stability
@@
Many LLMs struggle with limited stability: the response to a given input may vary widely depending on changes in input wording or semantics (newman2021), even as subtle as adding extra punctuation (formento2023). This not only presents a negative user experience to end users, requiring them to iterate on the same prompt (kim2024), but also open vulnerability to attacks that can lead to model output of harmful information (schulhoff2023). One common solution to this problem is model regularization, i.e. forcing the model to learn general patterns in the data, and methods for regularization include enforcing sparsity in model architecture (li2023), or more commonly augmenting training data with “noised” data points (chen2023). In the multimodal space, methods such as synonym replacement have shown to generate useful data points at inference time to reduce instability and improve overall performance (ge2023). We address the issue of stability through modality-grounded prompt perturbation and show that re-training on such prompts can help insulate models from performance drops on unseen prompts.
Methods
@@
We first outline the methods used for sampling data-grounded prompts for further model training (Prompt augmentation), identify the models trained using the augmented data (Models), and explain the QA data used in the experiments (Evaluation protocol).
Prompt augmentation
@@
Typical approaches to data augmentation include utilizing a variety of similar but distinct prompts to re-train the model, e.g. randomly changing the prompt instructions during training (lu2023). We test several simple methods for data augmentation to determine whether MFMs require modality-specific prompts or text-specific prompts to improve overall prompt stability.
@@
For a given text prompt $t$, we generate a list of $N=10$ perturbations $\hat{T}$ with a general-purpose LLM, using instructions to maximize diversity and semantic similarity to the original prompt. [^1] Llama-2, accessed 1 Jan 2024: https://huggingface.co/meta-llama/Llama-2-7b-chat-hf. We chose 10 as the number of perturbations to maximize prompt diversity without encouraging prompts that “drift” too far from the original prompt. This full set of perturbations $\hat{T}$ is used in evaluation to determine post-training performance of the models on the perturbed prompt data.
@@
We then compute the latent representation $X$ of text ($X_{t}$) and multimodal data ($X_{m}$) using the ImageBind model (girdhar2023), an encoder trained to align latent representations of data from different modalities with contrastive learning, using text as the intermediate modality. Finally, we select $k=3$ prompts from the set $\hat{T}$ that have maximum cosine similarity between the prompt text and the latent representations of the text, choosing either $X_{t}$ for text-similar paraphrases (“text-similar sampling”) or $X_{m}$ for modality-similar paraphrases (“modality-similar sampling”). We also sample randomly from the prompt perturbations as a control condition (“random sampling”). We choose 3 as the number of prompts per category to balance semantic diversity with amount of runtime available for training.
@@
In addition to these simpler conditions for prompt perturbation, we test a method to balance similarity of the perturbed prompt to the original data and diversity among the prompts. We sample 1 prompt perturbation $\hat{t}$ from $\hat{T}$ based on Eq. 1 similarity to original data (text and multimodal data).
[…]
@@
Then, in order to increase the diversity of the jointly sampled prompts, we normalize the similarity relative to the cumulative similarity of the original prompt to the set of already-sampled prompts ($\hat{T}^{\prime}$) from prior iterations (Eq. 2). As each additional prompt perturbation $\hat{t}$ is sampled, it is added to $\hat{T}^{\prime}$.
[…]
@@
This method (“Joint similarity”) yields a diverse set of input text that remains grounded with respect to the modality-specific data, thus decreasing the risk of re-training the model on augmented prompts that are irrelevant to the input context.
@@
Text similarity Modality similarity Joint-similarity
@@
Table 2: Prompt sample methods from Prompt augmentation, with example data and prompt (video frame from iVQA).
@@
For our experiments, we tested image, video, and spoken audio data [^2] At time of writing, no open-source generalist models proved capable of handling general audio in addition to other modalities, only human speech audio., which are three related but complementary data modalities that a generalist multimodal model should reasonably be expected to handle. Despite the proliferation of literature related to MFMs, relatively few models are available for open-source inference, and even fewer for fine-tuning on existing model weights. We investigate the following models which have released open-source implementations including training code:
@@
All the models encode multimodal data into the same latent representation space as text, provide text and multimodal data jointly to a universal processing unit, and generate output in an autoregressive fashion. We considered other models such as Macaw (lyu2023) and OneLLM (han2023) but did not adopt them due to the higher difficulty of implementation and seemingly poor performance outside of the image domain.
Evaluation protocol
@@
We evaluate our proposed method for data sampling using the following QA datasets that have free-response questions:
@@
  • VQA (antol2015): questions about MS COCO images
  • iVQA (yang2021): questions about how-to videos from YouTube
@@
  • Spoken-SQuAD (lee18): questions about spoken Wikipedia articles (generated via text-to-speech)
@@
We choose these tasks to approximate the kinds of tasks expected of MFMs, seen in the original work on OFASys and Unified-IO (lu2022; wang2022). We use a total of 1500 prompts and associated data (images, audio, video) for each dataset. Throughout evaluation (Results), we use the same 80%/20% train/test split of the data for fair comparison. Some of the models in this study had size limitations on the image and video data, which required extra data preprocessing: e.g., for images and videos, we pad and crop the data to 224 x 224 pixels; for videos, we sample 1-second frames for a total of 8 frames per video. We provide summary statistics about the data in Table TODO: summary statistics (description; question length; answer length; distribution of data size/length)
@@
To match the open-response format of the QA tasks, we evaluate the models’ performance using typical text generation metrics including BLEU, ROUGE, and BERT-Score (zhang2020). In general, we find that the BERT-Score results have higher values than BLEU and ROUGE, due to the fuzzy nature of the metric (semantic similarity) as compared to the more strict word-overlap metrics that penalize slight changes in wording.
Results
Testing on perturbed data
We first assess the degree to which multimodal models are able to correctly respond to perturbed prompts generated via three separate models:
  1. 1.
    General-purpose LLM (see Prompt augmentation)
  2. 2.
    Decoder model fine-tuned on paraphrasing [^5] Model accessed 30 Aug 2023: https://huggingface.co/ramsrigouthamg/t5˙sentence˙paraphraser.
  3. 3.
    Back-translation model (between English and Russian) [^6] Models accessed 30 Aug 2023: https://huggingface.co/Helsinki-NLP/opus-mt-en-ru https://huggingface.co/Helsinki-NLP/opus-mt-ru-en.
The results are shown in Table 3, in which the image modality has the most consistent drop in performance. We see consistent degradation of performance from the original questions to the perturbed prompts, up to a 50% drop in the case of Unified-IO and images (BLEU score of 0.687 for original prompts vs. 0.347 for LLM-paraphrase prompts). The video domain has slight degradation and low performance overall, due in part to the more complicated answers expected of the model (e.g., “The device being used in the video is a blender.”) as well as the more complex nature of videos in general. The overall low performance in the audio domain may be due to the inherently difficult nature of the QA task, which requires first converting speech to text and then performing logical reasoning over the text to answer the question. Surprisingly, both OFASys and Unified-IO reported high performance in audio and video tasks in the original work (e.g., up to 42.1% accuracy for OFASys on VQA for MSR-VTT data; bai2022 bai2022), which may indicate overfitting on a specific format of the task.
Based on this analysis, we conclude that prompt stability is a major concern for generalist multimodal models even in “strong” domains such as images, and a less substantial problem for less well-supported domains such as video.
@@ Results
Table 3: Performance of no-training model on original and perturbed prompts (standard error in parentheses).
Training on perturbed data
@@
We address the performance degradation on perturbed data by re-training the MFMs using the prompts generated through data augmentation (Prompt augmentation), applied to the QA datasets. We follow the same training procedure across all models, using the same hyperparameters7773 training epochs, learning rate $55e-$ 5, batch size 2, Adam optimizer with default hyperparameters. and objective function of cross-entropy on the model output logit probabilities compared to the ground-truth data. For the original prompt condition, we only train the model on the original prompts from all QA datasets. [^8] We also tried a condition where we trained on the same number of original prompts as the sample prompts, i.e. repeat each original prompt $N=3$ times in the training data, and found substantially worse performance across models and tasks. The decoder takes 99.3M parameters for OFASys and 409M parameters for Unified-IO, which is sufficiently lightweight to train on a single A100 GPU.
For each prompt perturbation condition, we use 3 sampled prompts from that condition in place of the original prompt (e.g. top-3 prompts with modality data similarity for “Modality-similarity” condition). OFASys and Unified-IO have a similar architecture that uses an LLM as the decoder for the combined input from text and other modalities, and we only retrain this decoder in both models as we are primarily interested in improving stability in the generated output, i.e. directly from the decoder, rather than improving e.g. encoder representation stability.
To summarize, we evaluate the following training conditions for each model type:
  • No training
  • Train only on original prompt (“Original”)
  • Train on prompt perturbations (“Joint similarity”, “Modality-similarity”, “Text similarity”, “Random sample”; see Prompt augmentation).
@@
We first show the model results on the original prompts in Table 4. The models without training (“No training”) perform poorly on audio and video and reasonably well on image. The models trained on the original prompts perform slightly better than the models trained on the perturbed prompts for the image domain, worse for the audio domain, and roughly the same for the video domain. The models trained on random samples of prompt perturbations generally perform worse than the other perturbation-trained models (e.g., image domain for OFASys), which may be due to noisy prompts that have little semantic relation with the meaning of the original prompt. In general, the results don’t show evidence of catastrophic forgetting, where the models re-trained on prompt perturbations would lose their ability to understand the original prompts due to overfitting on the perturbations.
@@
Table 4: Model accuracy on original prompts (standard error in parentheses).
Next, we evaluate model performance on all prompt perturbations ($\hat{T}$) generated for data augmentation (Prompt augmentation), to determine whether one method of prompt perturbation yields better generalization. The results are shown in Table 5. Across all modalities, the baseline model and the model trained on original prompts show substantial decrease in performance on prompt perturbations versus the original prompts. In contrast, the models trained on the prompt perturbations show consistently high performance on the prompt perturbations, with a slight decrease in the video modality and an increase in the audio modality. Surprisingly, the models show consistent performance regardless of the prompt sampling method used, which implies that simply exposing the model to a wider diversity of prompts always helps in the perturbation test setting, even if the exact form of the prompts varies. In the video domain, the regular prompt-trained model has substantially worse performance than the perturbation-trained models, possibly due to a more limited semantic space covered by the original prompts in the video domain as compared to the prompt perturbations (e.g. covering different verbs to describe the same motion or actions taken in the videos).
@@
Table 5: Model accuracy on prompt perturbations (standard error in parentheses).
@@
To assess the consistency of the performance differences over the prompt perturbations test data, we compute the coefficient of variation $CV$ ($\frac{\sigma^{2}}{\mu}$) over all the prompt perturbations, i.e. compute the standard deviation in performance score over all prompts for a given dataset/model and normalize by the mean. The normalized value represents the amount of variation in score relative to the expected score, e.g. a standard deviation of 5 in BLEU score for a high-scoring model should be better than the same standard deviation for a low-scoring model.
Shown in Table 6, the re-trained models achieve lower $CV$ on the prompt perturbations. This shows that training with any of the prompt perturbation strategies yields more stable performance as compared to training only on the original prompts. The variation is slightly higher in the audio domain, which aligns with the generally lower performance on all forms of prompt perturbations in audio (Table 3), such that even a small absolute error is relatively large in comparison to the overall lower performance.
@@
Modality Cluster theme Data description Prompt Perturb prompt True answer Joint similarity Original prompts Joint similarity Original prompts
@@
Table 7: Example data from target modalities, clustered by semantic representation. Cluster themes determined by manual inspection of prompts. Models organized by architecture and by training condition: “Original prompts” indicates models trained only original prompts, “Joint similarity” indicates models trained on prompt perturbations filtered for similarity to modality data and text data.
BLEU (mean)
Modality Cluster Example prompts Perturbation- training (all models) Original prompts Ratio
Audio 0 (religion) What do progressive moderates of Islam seek to separate? In the United States the UMC ranks as the largest what? What did the new mass allow as regards ceremony? 0.4755 0.0347 13.70
2 (space) What happened to the Apollo 13 crew to force them to use the LM as a ”life boat” as it had been designed? What is the name for O3 most often used? What did NASA name the series seeking up to 30 more flights to Earth’s orbit? 0.5583 0.0467 11.96
4 (law, history) What have the Treaties sought to enable since its foundation? Which courts have a duty to interpret domestic law as far as possible? How can you protest against the government in an individual way? 0.5300 0.0758 6.99
Image 3 (street views) What is on the sidewalk beside the woman? How many signs are on the pole? What is on the ladder rungs? 0.9050 0.3713 2.44
4 (people eating together) Where are these people eating? What color is the person’s shirt? Does this look more like a party or a funeral? 0.7344 0.3912 1.88
7 (Home activities) What color are the walls? What kind of shop is this? How many bags are black? 0.7049 0.3851 1.83
Video 2 (personal objects) What instrument is the person playing? What is the person wearing on her neck? What type chair is the person sitting on? 0.4131 0.0126 32.79
3 (hand movement) What is placed next to the bowl? What is she holding? What does the subject slide into the machine? 0.5262 0.0177 29.73
0 (personal grooming) What tool did she use? How was the girl combing her hair? What promenant fashion accessory is the lady wearing? 0.4195 0.0156 26.89
Table 8: Per-modality clusters and model scores, for perturbation-trained and prompt-trained OFASys models (other models omitted for space), sorted by ratio of BLEU score on prompt perturbations (mean for all perturbation-trained models, divided by prompt-trained model).
Error analysis

### v2

@@ Abstract
Surprisingly Fragile: Assessing and Addressing Prompt Instability in Multimodal Foundation Models
Ian Stewart, Sameera Horawalavithana, Brendan Kennedy, Sai Munikoti, Karl Pazdernik Pacific Northwest National Laboratory Richland, WA 99354 {ian.stewart,yasanka.horawalavithana,brendan.kennedy,sai.munikoti,karl.pazdernik}@pnnl.gov
1 Introduction
While text-only large language models (LLMs) are powerful in their own right, multimodal foundation models (MFMs) offer even more transformative capabilities such as image captioning and audio signal processing li2024. Generalist models such as Unified-IO are trained to process a range of modalities including image, video, and audio signal, often using an LLM as a combined encoder for all input types after preprocessing  lu2022. MFMs have the potential for major impact in downstream applications such as autonomous navigation cui2024 and medical image analysis moor2023, provided that their output is consistent in the face of variable user input.
Despite their power, these models often show surprising fragility in the face of prompts that have slight differences from the expected syntax or semantics seen during training newman2021. We show an example from a popular video QA dataset (iVQA) in Table 1, in which a person fixes a car engine and the model is asked to respond to the relevant prompt “What mechanism is the person working on?”. In response to prompt perturbations (paraphrased prompts), a baseline model with no additional training responds with a variety of incorrect responses. In contrast, a model that has been trained on prompt perturbations (“Joint similarity”; see Section 3.1) responds correctly regardless of phrasing.
@@ 1 Introduction
This study addresses a central research question in model robustness: to what degree does re-training multimodal models on prompt perturbations, chosen by different selection strategies, improve robustness? We investigate the relative stability of current state-of-the-art MFMs on several standard multimodal benchmark datasets and find the following:
@@
  • Modern MFMs show a severe drop in performance on the image domain, in response to perturbed prompts on standard QA datasets (4.1).
@@
  • Re-training MFMs on a wider variety of perturbed prompts, regardless of how the prompts are selected (3.1), improves performance across domains and reduces instability on other prompt perturbations, more than simply re-training on the original prompts (4.2).
  • The perturbation-trained models show consistent improvement when the prompts target modality-specific information, e.g. prompts in the image modality that reference material properties (4.3).
@@
Our work provides an automated method to help align MFMs to human expectations, such as consistency in response to input variation zhao2024.
2 Related work
2.1 Multimodal foundation models
@@
Recent work has unified text-based foundation models with modality-specific processing to yield MFMs that can respond to input from a variety of modalities li2024. The majority of such models, including Unified-IO and OFASys, first encode the text and multimodal data to a shared representation space, then decode the combined representation into the desired modality, often text (e.g. image QA requires encoding image + text data and decoding as a text response). Some systems such as Next-GPT learn to perform any-to-any modality generation by training modality-specific decoders wu2023, which works reasonably well for more common modalities such as images. For breadth, systems including ImageBind girdhar2023 and OneLLM han2023 use a universal encoder and learn from a wide variety of data including depth maps and point clouds, which may help the model’s internal LLM achieve better cross-modality understanding. To further improve performance, parameter-efficient methods such as prompt learning have helped to solve problems with missing data lee2023 and cross-modality alignment khattak2023. We identify a possible gap in the research space, i.e. model-agnostic methods to improve robustness, and test a variety of prompt perturbation methods to determine the relative impact on task performance.
2.2 Model stability
@@
Many LLMs struggle with limited stability: the response to a given input may vary widely depending on changes in input wording or semantics newman2021, even as subtle as adding extra punctuation formento2023. This not only presents a negative user experience to end users, requiring them to iterate on the same prompt kim2024, but also opens vulnerability to attacks that can lead to model output of harmful information schulhoff2023. One common solution to this problem is model regularization, i.e. forcing the model to learn general patterns in the data, and methods for regularization include enforcing sparsity in model architecture li2023, or more commonly augmenting training data with “noised” data points chen2023. In the multimodal space, methods such as synonym replacement have shown to generate useful data points at inference time to reduce instability and improve overall performance ge2023. We address the issue of stability through modality-grounded prompt perturbation and show that re-training on prompts, regardless of sampling strategy, can help protect models from performance drops on unseen prompts.
3 Methods
@@
We first outline the methods used for sampling data-grounded prompts for further model training (3.1), identify the models trained using the augmented data (3.2), and explain the QA data used in the experiments (3.3).
3.1 Prompt augmentation
@@
Typical approaches to data augmentation include utilizing a variety of similar but distinct prompts to re-train the model, e.g. randomly changing the prompt instructions during training lu2023. We test several simple methods for data augmentation to determine whether MFMs require modality-specific prompts or text-specific prompts to improve overall prompt stability.
@@
For a given text prompt $t$, we generate a list of $N=10$ perturbations $\hat{T}$ with a general-purpose LLM, [^1] Llama-2, accessed 1 Jan 2024: https://huggingface.co/meta-llama/Llama-2-7b-chat-hf. using instructions to maximize diversity and semantic similarity to the original prompt. [^2] Instruction: “Write 10 new questions that are paraphrases of the first question. The new questions should be unique but should have the same meaning as the first question. First question:” We choose 10 as the number of perturbations to maximize prompt diversity while avoiding prompts that “drift” too far from the original prompt (e.g., generating 100 perturbations could yield highly divergent perturbations). This full set of perturbations $\hat{T}$ is used in evaluation to determine post-training performance of the models on the perturbed prompt data.
@@
We then compute the latent representation $X$ of text ($X_{t}$) and multimodal data ($X_{m}$) using the ImageBind model girdhar2023, an encoder trained to align latent representations of data from different modalities with contrastive learning, using text as the intermediate modality. Finally, we select $k=3$ prompt perturbations from the set $\hat{T}$ with maximum cosine similarity between the prompt perturbation text and the data $X$. In this step, we choose either $X_{t}$ for perturbations most similar to the text data (“text-similar sampling”) or $X_{m}$ for perturbations most similar to the modality data (“modality-similar sampling”). As a control condition, we also sample randomly from the prompt perturbations (“random sampling”). We choose 3 as the number of prompts per category to balance semantic diversity with a reasonable runtime for training and a need to avoid overfitting.
@@
In addition to these simpler conditions for prompt perturbation, we test a method to balance similarity of the perturbed prompt to the original data and diversity among the prompts. In the first step, we sample 1 prompt perturbation $\hat{t}$ from $\hat{T}$ based on Eq. 1 similarity to original data (text and multimodal data).
$P(\hat{t})\propto\text{JointSim}(\hat{t})=sim(\hat{t},X_{t})+sim(\hat{t},X_{m})$ (1)
@@
For all subsequent prompt samples, in order to increase the diversity of the jointly sampled prompts, we normalize the similarity relative to the cumulative similarity of the original prompt to the set of already-sampled prompts ($\hat{T}^{\prime}$) from prior iterations (Eq. 2). As each additional prompt perturbation $\hat{t}$ is sampled, it is added to $\hat{T}^{\prime}$.
$P(\hat{t})\propto\frac{\text{JointSim}(\hat{t})}{\frac{1}{|\hat{T}^{\prime}|}\sum_{\hat{t}^{\prime}\in\hat{T}^{\prime}}sim(\hat{t},\hat{t}^{\prime})}$ (2)
@@
This method (“Joint similarity”) yields prompt perturbations that are divergent from one another and also grounded with respect to both text data $X_{t}$ and modality data $X_{m}$, thus decreasing the risk of re-training the model on augmented prompts that are irrelevant to the input context.
@@
Text similarity Modality similarity Joint similarity
@@
Table 2: Prompt perturbation sampling methods from 3.1, with example data and prompt (video frame from iVQA).
@@ 3.2 Models
For our experiments, we tested image, video, and spoken audio data, [^3] At time of writing, no open-source generalist models prove capable of using general-domain audio as input, only human speech audio. which are three related but complementary data modalities that a MFM should reasonably be expected to handle. Despite the proliferation of literature related to MFMs, relatively few models are available for open-source inference, and even fewer for fine-tuning on new datasets without extensive preprocessing. We investigate the following models which have released open-source implementations including training code:
@@
All the models encode multimodal data into the same latent representation space as text, pass text and multimodal data jointly to a universal processing unit, and generate output in an autoregressive fashion. We considered other models such as Macaw lyu2023 and OneLLM han2023 but did not adopt them due to the higher difficulty of implementation and seemingly poor performance outside of the image domain, even after fine-tuning.
3.3 Evaluation protocol
@@
We evaluate the proposed methods for data sampling using the following QA datasets that have free-response questions:
@@
  • VQA antol2015: questions about MS COCO images depicting everyday life scenes;
  • iVQA yang2021: questions about how-to videos from YouTube;
@@
  • Spoken-SQuAD lee18: questions about spoken Wikipedia articles (audio via text-to-speech).
@@
We choose these tasks to approximate the kinds of tasks expected of MFMs, seen in the original work on OFASys and Unified-IO lu2022; wang2022. We use a total of 1500 prompts and associated data (images, audio, video) for each dataset. Throughout evaluation (4), we use the same 80%/20% train/test split of the data for fair comparison. Some of the models in this study have size limitations on the image and video data, which require extra data preprocessing. For example: for images and videos, we pad and sub-sample the data to 224 x 224 pixels; for videos, we sample 1-second frames for a total of 8 frames per video.
@@
To match the open-response format of the QA tasks, we evaluate the models’ performance using typical text generation metrics including BLEU, ROUGE, and BERT-Score zhang2020. In general, we find that the BERT-Score results have higher values than BLEU and ROUGE, due to the fuzzy nature of the metric (semantic similarity) as compared to the more strict word-overlap metrics that penalize slight changes in wording.
@@
Table 3: Performance of baseline models (no additional training) on original and perturbed prompts; standard error in parentheses.
4 Results
4.1 Testing on perturbed data
We first assess the accuracy of the baseline MFMs in response to perturbed prompts generated via three separate models:
  1. 1.
    General-purpose LLM (see 3.1)
  2. 2.
    Decoder model fine-tuned on paraphrasing [^6] Model accessed 30 Aug 2023: https://huggingface.co/ramsrigouthamg/t5_sentence_paraphraser.
  3. 3.
    Back-translation model (between English and Russian; chosen for sufficient linguistic distance to increase distance from original prompt) [^7] Models accessed 30 Aug 2023: https://huggingface.co/Helsinki-NLP/opus-mt-en-ru https://huggingface.co/Helsinki-NLP/opus-mt-ru-en.
The results are shown in Table 3, in which the image modality has the most consistent drop in performance. We see consistent degradation of performance from the original questions to the perturbed prompts, up to a 50% drop in the case of Unified-IO and images (BLEU score of 0.687 for original prompts vs. 0.347 for LLM-paraphrase prompts). The video domain has slight degradation and low performance overall, due in part to the more complicated answers expected of the model (e.g., “The device being used in the video is a blender.”) as well as the more complex nature of videos in general. The overall low performance in the audio domain may be due to the inherently difficult nature of the QA task, which requires first converting speech to text and then performing logical reasoning over the text to answer the question. Surprisingly, both OFASys and Unified-IO reported high performance in audio and video tasks in the original work (e.g., up to 42.1% accuracy for OFASys on VQA for MSR-VTT data; bai2022 bai2022), which may indicate overfitting on a specific format of the task. We provide additional results in appendix˜A using modern models (Phi-4 and Gemma-3n), which support the finding of performance degradation on text data and inconsistent trends in modality-specific data perturbations.
Based on this analysis, we conclude that prompt stability is a major concern for generalist multimodal models even in “strong” domains such as images, and a more minor problem for less well-supported domains such as video.
4.2 Training on perturbed data
@@
We address the performance degradation on perturbed data by re-training the MFMs using the prompts generated through data augmentation (3.1), applied to the QA datasets. We follow the same training procedure across all models, using the same hyperparameters8883 training epochs, learning rate 5e55e-5, batch size 2, Adam optimizer with default hyperparameters. and objective function of cross-entropy on the model output logit probabilities compared to the ground-truth data. For the original prompt condition, we only train the model on the original prompts from all QA datasets. [^9] We also tried a condition where we trained on the same number of original prompts as the sample prompts, i.e. repeat each original prompt $N=3$ times in the training data, and found substantially worse performance across models and tasks. The decoder takes 99.3M parameters for OFASys and 409M parameters for Unified-IO, which is sufficiently lightweight to train on a single A100 GPU.
For each prompt perturbation condition, we use 3 sampled prompts from that condition in place of the original prompt: e.g., for the “Modality-similarity” condition, we use the top-3 prompt perturbations based on similarity to the modality data. OFASys and Unified-IO have a similar architecture that uses an LLM as the decoder for the combined input from text and other modalities, and we only retrain this decoder in both models as we are primarily interested in improving stability in the generated output directly from the decoder, rather than improving stability in e.g. encoder representation.
To summarize, we evaluate the following training conditions for each model type: No training; Train only on original prompt (“Original”); Train on prompt perturbations (“Joint similarity”, “Modality-similarity”, “Text similarity”, “Random sample”; see 3.1).
@@ Results
We first show the model results on the original prompts in Table 4. The models without training (“No training”) perform poorly on audio and video and reasonably well on image. The models trained on the original prompts perform slightly better than the models trained on the perturbed prompts for the image domain, worse for the audio domain, and roughly the same for the video domain. The models trained on random samples of prompt perturbations generally perform worse than the other perturbation-trained models (e.g., image domain for OFASys), which may be due to noisy prompts that have little semantic relation with the meaning of the original prompt. In general, the results don’t show evidence of catastrophic forgetting kemker2018, where the models re-trained on prompt perturbations completely lose the ability to understand the original prompts due to overfitting on the perturbations.
@@
Table 4: Model accuracy on original prompts; standard error in parentheses.
Next, we evaluate model performance on all prompt perturbations ($\hat{T}$) generated for data augmentation (3.1), to determine whether a particular form of data-grounded prompt perturbation yield better generalization. The results are shown in Table 5. Across all modalities, the baseline model and the model trained on original prompts show substantial decrease in performance on prompt perturbations versus the original prompts. In contrast, the models trained on the prompt perturbations show consistently high performance on the prompt perturbations, with a slight decrease in the video modality and an increase in the audio modality. In the video domain, the original prompt-trained model has substantially worse performance than the perturbation-trained models, possibly due to a more limited semantic space in the original prompts in the video domain as compared to the prompt perturbations (e.g. perturbations contain variety of verbs to describe the same motion or actions taken in the videos).
Surprisingly, the perturbation-trained models show consistent performance regardless of the prompt sampling method used. This implies that simply exposing the model to a wider diversity of prompts always helps when testing on unseen perturbations, even if the exact nature of the generated prompts isn’t consistent.
@@
Table 5: Model accuracy on prompt perturbations; standard error in parentheses.
@@
To assess the consistency of the performance differences over the prompt perturbations test data, we compute the coefficient of variation CVCV ($\frac{\sigma}{\mu}$) over all the prompt perturbations. Using the performance scores for all output from a given model and task, we compute the standard deviation in scores, and normalize the standard deviation of the scores by the mean. The normalized value represents the amount of variation in score relative to the expected score, e.g. a standard deviation of 5 in BLEU score for a high-scoring model should be better than the same standard deviation for a low-scoring model.
Shown in Table 6, the re-trained models achieve lower $CV$ on the prompt perturbations. This again confirms that training with any of the prompt perturbation strategies yields more stable performance as compared to training only on the original prompts. The variation is slightly higher in the audio domain, which aligns with the generally lower performance on all forms of prompt perturbations in audio (Table 3), such that even a small absolute error is relatively large in comparison to the overall lower performance.
@@
Modality Cluster theme Example data Prompt Perturb prompt True answer Joint similarity Original prompts Joint similarity Original prompts
@@
Table 7: Example data from target modalities, clustered by semantic representation. Cluster themes determined by manual inspection of prompts. Models organized by architecture and by training condition: “Original prompts” indicates the model trained only with original prompts, “Joint similarity” indicates the model trained on prompt perturbations filtered for similarity to modality data and text data.
4.3 Error analysis